# A Catalog of First-Order Electroweak Phase Transitions in the Standard Model Effective Field Theory

### v1

@@ Appendix A Simplified analysis of Higgs mass and vacuum expectation value
$J_{a}(x)$ $=\frac{x^{2}}{64\pi^{2}}\left(\log\frac{x}{\mu^{2}}-C_{a}\right),\quad C_{a}=\begin{cases}\frac{3}{2}&\text{for scalars and fermions,}\\ \frac{5}{6}&\text{for gauge bosons,}\\ \end{cases}$ (A.3)
$J_{a}^{T\neq 0}(x)$ $=\int_{0}^{\infty}\mathrm{d}{y}~{}y^{2}\log\left(1\pm e^{-\sqrt{y^{2}+x}}\right),\quad\begin{cases}+&\text{for fermions,}\\ -&\text{for bosons.}\\ \end{cases}$ (A.4)
@@ Appendix B Estimating coefficients with cancellations
where we replaced $-C^{\varphi}T^{2}$ with $C^{\varphi}_{3}$, which holds at leading order. Assuming that $C^{\varphi}_{3}$ is never expected to be larger (in a power-counting sense) than $\lambda$, we can hence use the size of $\lambda$ to estimate $\lambda_{3}/T$. Unless there is a cancellation between the two terms, of course.
@@ C.1 Tree-level terms only
    $-1<n_{\phi}<1,$ (C.14)
    from which we can pick $n_{\phi}=0$ as a representative. This is the case “TLB” in the catalog, and the counting is
@@
    $-1<n_{\phi}<0,$ (C.16)
    from which we can pick $n_{\phi}=-1/2$. This is the case “SQTLB” in the catalog, and the counting is
@@ C.2 Supercooled tree-level barriers
[…]
The simple difference between the regular TLB case and the supercooled variant considered here, is in the strict equality of (C.22). This opens up the possibility space further. First of all, it is possible to have a larger $n_{C}$ than for TLB and SQTLB, and the rest of the parameters the same. This automatically satisfies the power counting restrictions, and will have the same properties in general as the regular potentials. We label these supercooled cases “SC TLB (1)”, with SC short for supercooled, and similarly for SQTLB.
@@
    from which we can pick $n_{\phi}=1/2$ as a representative. This is what we label as “SC TLB (2)” in the catalog, and the counting is
@@
    $-1/2<n_{\phi}<0,$ (C.30)
    from which we can pick $n_{\phi}=-1/4$. This is the case “SC SQTLB (2)” in the catalog, and the counting is
@@ C.3 Radiative barrier
This counting is analogous to the simple example we studied in section 3, with the caveat that here also the $\phi^{6}$ term is included. But because the $\phi^{6}$ coefficient did not play a role in the derivation above, we can also use the same counting but with $n_{C}>3$ to represent potentials with radiative barriers but no $\phi^{6}$ term. Hence, we have here derived the countings named “RB” and “RB 6” in the catalog.
@@ C.4 Radiative symmetry breaking
[…]
@@
This counting models radiative symmetry breaking. Because the $\phi^{6}$ coefficient did not play a role in the derivation above, we can also use the same counting but with $n_{C}>6$ to represent potentials with radiative symmetry breaking but no $\phi^{6}$ term. Hence, we have here derived the countings named “RSB” and “RSB 6” in the catalog.
@@ C.5 Supercooled radiative barrier and radiative symmetry breaking
    from which we can pick $n_{\phi}=1/2$. This is the case of “supercooled RB” in the catalog, and the counting is [^9] The mass scale $\left(g^{\frac{7}{2}}T^{2}\right)^{\frac{1}{2}}=g^{\frac{7}{4}}T$ lies between the supersoft and the ultrasoft scales—see (2.3).
@@
    $-1<n_{\phi}<1,$ (C.62)
    from which we can pick $n_{\phi}=0$. This is the case of “supercooled RSB” in the catalog, and the counting is

### v2

@@ Appendix A Simplified analysis of Higgs mass and vacuum expectation value
[…]
@@ Appendix B Estimating coefficients with cancellations
where we replaced $C^{\varphi}T^{2}$ with $C^{\varphi}_{3}$, which holds at leading order. Assuming that $C^{\varphi}_{3}$ is never expected to be larger (in a power-counting sense) than $\lambda$, we can hence use the size of $\lambda$ to estimate $\lambda_{3}/T$. Unless there is a cancellation between the two terms, of course.
@@ C.1 Tree-level terms only
    $1<n_{\phi}<1,$ (C.14)
    from which we can pick $n_{\phi}=0$ as a representative. This is the case ”TLB” in the catalog, and the counting is
@@
    $1<n_{\phi}<0,$ (C.16)
    from which we can pick $n_{\phi}=-1/2$. This is the case ”SQTLB” in the catalog, and the counting is
@@ C.2 Supercooled tree-level barriers
(C.23) $\implies n_{\phi}<3-n_{\lambda}.$ (C.27)
The simple difference between the regular TLB case and the supercooled variant considered here, is in the strict equality of (C.22). This opens up the possibility space further. First of all, it is possible to have a larger $n_{C}$ than for TLB and SQTLB, and the rest of the parameters the same. This automatically satisfies the power counting restrictions, and will have the same properties in general as the regular potentials. We label these supercooled cases ”SC TLB (1)”, with SC short for supercooled, and similarly for SQTLB.
@@
    from which we can pick $n_{\phi}=1/2$ as a representative. This is what we label as ”SC TLB (2)” in the catalog, and the counting is
@@
    $1/2<n_{\phi}<0,$ (C.30)
    from which we can pick $n_{\phi}=-1/4$. This is the case ”SC SQTLB (2)” in the catalog, and the counting is
@@ C.3 Radiative barrier
This counting is analogous to the simple example we studied in section 3, with the caveat that here also the $\phi^{6}$ term is included. But because the $\phi^{6}$ coefficient did not play a role in the derivation above, we can also use the same counting but with $n_{C}>3$ to represent potentials with radiative barriers but no $\phi^{6}$ term. Hence, we have here derived the countings named ”RB” and ”RB 6” in the catalog.
@@ C.4 Radiative symmetry breaking
$\mathbf{A}:\quad n_{m^{2}}=2$ $<8\quad(\text{OK!}),$ (C.50)
$\mathbf{B}:\quad n_{m^{2}}=2$ $>0\quad(\text{OK!}),$ (C.51)
$\mathbf{(C)}:\quad n_{m^{2}}=2$ $>0\quad(\text{OK!}).$ (C.52)
@@
This counting models radiative symmetry breaking. Because the $\phi^{6}$ coefficient did not play a role in the derivation above, we can also use the same counting but with $n_{C}>6$ to represent potentials with radiative symmetry breaking but no $\phi^{6}$ term. Hence, we have here derived the countings named ”RSB” and ”RSB 6” in the catalog.
@@ C.5 Supercooled radiative barrier and radiative symmetry breaking
    from which we can pick $n_{\phi}=1/2$. This is the case of ”supercooled RB” in the catalog, and the counting is [^10] The mass scale $\left(g^{\frac{7}{2}}T^{2}\right)^{\frac{1}{2}}=g^{\frac{7}{4}}T$ lies between the supersoft and the ultrasoft scales—see (2.3).
@@
    $1<n_{\phi}<1,$ (C.62)
    from which we can pick $n_{\phi}=0$. This is the case of ”supercooled RSB” in the catalog, and the counting is